OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2

OpenStudy (anonymous):

x=4

OpenStudy (ash2326):

3x+2y=8 2x+3y=2 multiply 1st by 3 we get 9x+6y=24 multiply 2nd equation by 2 4x+6y=4 subtract the second from first 5x=20 x=4 put x in first equation 12+2y=8 2y=-4 y=-2 x=4 and y=-2